OpenStudy (anonymous):

A round 4 ft wide hot tub is being built into a 16 ft by 26 ft rectangular deck. How much decking is needed for the surface of the deck?

OpenStudy (anonymous):

\[\pi\]

OpenStudy (anonymous):

If I understood the question correct, you need to subtract the area of the hot tub from the area of the rectangular deck. If the hot tub is 4 ft wide, then the radius is 2 ft. For the circle A=πr^2, for the rectangular A=l*w A=πr^2=3.14*2^2=12.56 A=l*w=26*16=416 416-12.56=403.44